# Classification of fractional, singular Yamabe metrics on a twice-punctured sphere I

### v1

@@ 6.2.1. Local behavior near isolated singularities
Using a blow-up analysis, Caffarelli, Jin, Sire and Xiong [MR3366748] show that any isolated singularity of a singular Yamabe metric on the sphere must be asymptotically radial. Combining this with our classification of the radial solutions, we see that any isolated singularity is asymptotic to a Delaunay solution. In other words, if $s<1$ is sufficiently close to 11 and if $\Lambda\subset\mathbf{S}^{n}$ is a finite set and $U:\mathbf{S}^{n}\backslash\Lambda\rightarrow(0,\infty)$ satisfies the boundary condition and solves the PDE
@@
In the case that $1-\delta<s<1$ with $\delta>0$ sufficiently small, we can use the slide-back techniques of [MR1666838] to strengthen this estimate, showing there exist $T\in\mathbf{R}$, $a\in\mathbf{R}^{n}$, and $\beta>1$ such that
@@
Here, once again, the dot denotes a derivative with respect to $t$. We expect that this proof should be very similar to the one in [MR1666838], so we provide the strategy for the proof below and leave its details for subsequent work.
We start with a solution
$v:[0,\infty)\times\mathbf{S}^{n-1}\rightarrow(0,\infty),\quad P_{s}v+c_{n,s}v=c_{n,s}v^{\frac{n+2s}{n-2s}}$
and aim to show that there exists a Delaunay asymptote $v_{\varepsilon}$ and parameters $T\in\mathbf{R}$, $a\in\mathbf{R}^{n}$, and $\beta>1$ such that the estimate (43) holds. Let $\{\tau_{k}\}_{k\in\mathbf{N}}\subset\mathbf{R}$ be any sequence of real numbers such that $\tau_{k}\to\infty$ as $k\to\infty$ and define the so-called slide-back sequence $v_{k}:[-\tau_{k},\infty)\times\mathbf{S}^{n-1}\rightarrow(0,\infty)$ as
@@
By the a priori estimates proven in § 3.2 , we can use the Arzela-Ascoli theorem to extract a convergent subsequence $v_{k}\rightarrow v_{*}$, which (by our classification result) must have the form
@@
To obtain simple asymptotics we must show $\varepsilon$ and $T$ do not depend of the choice of the sequence $\{\tau_{k}\}_{k\in\mathbf{N}}\subset\mathbf{R}$. We prove the independence of $\varepsilon$ using the radial Pohozaev invariant (see (44) below); here it remains for us to prove that different Delaunay solutions in fact have different Pohozaev invariants. Next we prove the independence of $T$ follows from a delicate rescaling argument similar to the proof of [MR1666838, Proposition 5].
Once we have simple asymptotics, we use the nondegeneracy sketched in the previous subsection together with a version of the linear decomposition lemma (see [MR1356375, Lemma 4.18]) to get the refined asymptotics.
@@ 6.2.2. Compactness of the moduli space of complete metrics
As a final application, we describe a compactness result modeled on Pollack’s compactness theorem [MR1266101] in the scalar curvature setting. To state the result properly, we define both the marked moduli space and the unmarked moduli space of singular Yamabe metrics. If $\Lambda\subset\mathbf{S}^{n}$ is finite, we let
@@
The compactness result we prove is the following: if $\Omega\subset\mathcal{M}_{k}$ is such that the radial Pohozaev invariants of all Delaunay asymptotes are bounded away from zero and the distance between singular points remains bounded away from zero, then any sequence in $\Omega$ has a convergent subsequence. Again, we expect the proof to be similar in spirit to that of [MR1266101], [MR4504918], and [cmpt_6th_order], so we provide only the central ideas of the proof.
@@
Since it plays such a central role, we first take some time to define the (radial) Pohozaev invariant as (see [Azahara2018, Theorem 1.2]) [^1] Please notice that there is a typo in the coefficients $e_{1}^{*}$ and $e^{*}$ as written in [Azahara2018], but the asymptotic behavior is correct.
$\mathcal{P}_{s}(v)=\frac{c_{n,s}}{\widetilde{d}_{s}}\left(\frac{n-2s}{2n}v^{\frac{2n}{n-2s}}-\frac{1}{2}v^{2}\right)+\frac{1}{2}\int_{0}^{\rho_{0}^{*}}(\rho^{*})^{1-2s}\left(e_{1}^{*}(\rho)(\partial_{t}V)^{2}-e^{*}(\rho)(\partial_{\rho^{*}}V)^{2}\right)d\rho^{*},$ (44)
@@
where $V\in\mathcal{C}^{\infty}(\mathbf{R}^{2}_{+})$ is the Caffarelli-Silvestre extension of $v\in\mathcal{C}^{\infty}(\mathbf{R})$ and $e_{1}^{*}(\rho)$, $e_{*}(\rho)$ are positive coefficients depending only on $\rho$ and the dimention and such as $e_{1}^{*}(\rho),e_{*}(\rho)\to 1,$ as $\rho\to 0$. Here we recall that in Emden–Fowler coordinates, the bulk (Poincaré‐-Einstein) hyperbolic metric takes the form
$g_{+}=d\rho^{*}+\bigl(1+\tfrac{1-s}{2+2s}(\rho^{*})^{n-1}\bigr)dt^{2}+\mathcal{O}(\rho^{*})dtd\rho^{*},$
@@
where $\rho^{*}\in(0,\rho_{0}^{*})$ with $\rho_{0}^{*}>0$ being the critical extension parameter given by [Azahara2018, Corollary 4.2]. We emphasize that the formula above holds only for the conformal factor expressed in log-cylindrical coordinates. For the spherical and Euclidean gauge, we also set
$\mathcal{P}_{s}(u)=(\mathcal{P}_{s}\circ\mathfrak{F}^{-1})(u)\quad{\rm and}\quad\mathcal{P}_{s}(U)=(\mathcal{P}_{s}\circ\mathfrak{F}^{-1})(u\circ\Pi),$
@@
where $v\in\mathcal{C}^{\infty}(\mathbf{R})$, $u\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$, and $U\in\mathcal{C}^{\infty}(\mathbf{S}^{n}\setminus\{N,S\})$ are solutions to ($\widehat{\mathcal{Q}}_{n,s}$), ($\overline{\mathcal{Q}}_{n,s}$), and (${\mathcal{Q}}^{\circ}_{n,s}$), respectively. Here we recall that $\Pi:\mathbf{S}\setminus\{N,S\}\to\mathbf{R}^{n}\setminus\{0\}$ is the stereographic projection and $\mathfrak{F}:\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})\to\mathcal{C}^{\infty}(\mathbf{R}\times\mathbf{S}^{n-1})$ is the log-cylindrical transform. Let $\{g_{k}=U_{k}^{\frac{4}{n-2s}}g_{\circ}\}$ be a sequence of singular Yamabe metrics such that all the asymptotic radial Pohozaev invariants are bounded away from zero, and the distances between singular points are also bounded away from zero. Using (for instance) [MR4504918, Lemma 11], it suffices to assume the singular set is a fixed finite set $\Lambda$ for the entire sequence. Now, let $\{D_{\ell}\}_{\ell\in\mathbf{N}}$ be a compact exhaustion of $\mathbf{S}^{n}\backslash\Lambda$ and for any large $\ell\gg 1$, we extract a convergent subsequence $U_{k}\rightarrow U_{*}(\ell)$. Letting $\ell\rightarrow\infty$, we obtain a limit $U_{*}\in\mathcal{C}^{2,\alpha}(\mathbf{S}^{n}\backslash\Lambda)$ for some $\alpha\in(0,1)$. Finally, we use the fact that the Pohozaev invariants are bounded away from zero to show that $U_{*}>0$ is strictly positive and defines a complete metric on $\mathbf{S}^{n}\backslash\Lambda$, i.e. $\lim_{p\to\Lambda}U(p)=\infty$. If $U_{*}\equiv 0$, we could rescale the sequence and contradict the positive lower bound for all Pohozaev invariants.
@@
In the regime in which we’ve established our classification, namely $1-\delta<s<1$, we expect the proof described above to carry over from the local setting to the nonlocal setting without any complications. It is worth remarking that in the local setting one can write the radial Pohozaev invariant very explicitly in terms of the necksize, while such a relation in the nonlocal setting is more elusive, mostly because the expression (44) is much more complicated. Thus it would be very interesting to relate the necksize more explicitly to the radial Pohozaev invariant in the nonlocal setting.

### v2

@@ 6.2.1. Local behavior near isolated singularities
Using a blow-up analysis, Caffarelli, Jin, Sire and Xiong [6] show that any isolated singularity of a singular Yamabe metric on the sphere must be asymptotically radial. Combining this with our classification of the radial solutions, we see that any isolated singularity is asymptotic to a Delaunay solution. In other words, if $s<1$ is sufficiently close to 11 and if $\Lambda\subset\mathbf{S}^{n}$ is a finite set and $U:\mathbf{S}^{n}\backslash\Lambda\rightarrow(0,\infty)$ satisfies the boundary condition and solves the PDE
@@
In the case that $1-\delta<s<1$ with $\delta>0$ sufficiently small, we can use the slide-back techniques of [24] to strengthen this estimate, showing there exist $T\in\mathbf{R}$, $a\in\mathbf{R}^{n}$, and $\beta>1$ such that
@@
Here, once again, the dot denotes a derivative with respect to $t$. We expect that this proof should be very similar to the one in [24], so we provide the strategy for the proof below and leave its details for subsequent work.
We start with $v\in\mathcal{C}^{\infty}([0,\infty)\times\mathbf{S}^{n-1})$ a positive smooth solution to
$P_{s}v+c_{n,s}v=c_{n,s}v^{\frac{n+2s}{n-2s}}\quad{\rm in}\quad[0,\infty)\times\mathbf{S}^{n-1}$
and aim to show that there exists a Delaunay asymptote $v_{\varepsilon}$ and parameters $T\in\mathbf{R}$, $a\in\mathbf{R}^{n}$, and $\beta>1$ such that the estimate (44) holds. Let $\{\tau_{k}\}_{k\in\mathbf{N}}\subset\mathbf{R}$ be any sequence of real numbers such that $\tau_{k}\to\infty$ as $k\to\infty$ and define the so-called slide-back sequence $\{v_{k}\}_{k\in\mathbf{N}}\subset\mathcal{C}^{2}([-\tau_{k},\infty)\times\mathbf{S}^{n-1})$ as
@@
By the a priori estimates proven in Section 3.2 , we can use the Arzela-Ascoli theorem to extract a convergent subsequence $v_{k}\rightarrow v_{*}$, which (by our classification result) must have the form
@@
To obtain simple asymptotics, we must show that both $\varepsilon,T>0$ do not depend on the choice of the sequence. We prove the independence of $\varepsilon$ using the radial Pohozaev invariant (Hamiltonian) in (45) below; here it remains for us to prove that different Delaunay solutions in fact have different Pohozaev invariants. Next we prove the independence of $T$ follows from a delicate rescaling argument similar to the proof of [24, Proposition 5].
Once we have simple asymptotics, we use the nondegeneracy sketched in the previous subsection together with a version of the linear decomposition lemma (see [27, Lemma 4.18]) to get the refined asymptotics.
@@ 6.2.2. Compactness of the moduli space of complete metrics
As a final application, we describe a compactness result modeled on Pollack’s compactness theorem [29] in the scalar curvature setting. To state the result properly, we define both the marked moduli space and the unmarked moduli space of singular Yamabe metrics. If $\Lambda\subset\mathbf{S}^{n}$ is finite, we let
@@
The compactness result we prove is the following: if $\Omega\subset\mathcal{M}_{k}$ is such that the radial Pohozaev invariants (Hamiltonian) of all Delaunay asymptotes are bounded away from zero and the distance between singular points remains bounded away from zero, then any sequence in $\Omega$ has a convergent subsequence. Again, we expect the proof to be similar in spirit to that of [29], [2], and [3], so we provide only the central ideas of the proof.
@@
Since it plays such a central role, we first take some time to define the (radial) Pohozaev invariant as (see [17, Theorem 1.2]) [^1] Please notice that there is a typo in the coefficients $e_{1}^{*}$ and $e^{*}$ as written in [17], but the asymptotic behavior is correct.
$\mathcal{H}^{*}_{s}(v)=\frac{c_{n,s}}{\widetilde{d}_{s}}\left(\frac{n-2s}{2n}v^{\frac{2n}{n-2s}}-\frac{1}{2}v^{2}\right)+\frac{1}{2}\int_{0}^{\rho_{0}^{*}}(\rho^{*})^{1-2s}\left(e_{1}^{*}(\rho)(\partial_{t}V)^{2}-e^{*}(\rho)(\partial_{\rho^{*}}V)^{2}\right)d\rho^{*},$ (45)
@@
where $V\in\mathcal{C}^{\infty}(\mathbf{R}^{2}_{+})$ is the Caffarelli-Silvestre extension of $v\in\mathcal{C}^{\infty}(\mathbf{R})$ and $e_{1}^{*},e^{*}\in\mathcal{C}^{\infty}(0,\rho_{0}^{*})$ are positive coefficients depending only on $\rho$ and the dimension such that $e_{1}^{*}(\rho),e_{*}(\rho)\to 1$ as $\rho\to 0$, where $\rho^{*}\in(0,\rho_{0}^{*})$ with $\rho_{0}^{*}>0$ being the critical extension parameter given by [17, Corollary 4.2]. Here we recall that in Emden–Fowler coordinates, the bulk (Poincaré‐-Einstein) hyperbolic metric takes the form
$g_{+}=d\rho^{*}+\bigl(1+\tfrac{1-s}{2+2s}(\rho^{*})^{n-1}\bigr)dt^{2}+\mathcal{O}(\rho^{*})dtd\rho^{*}.$
@@
We emphasize that the formula above holds only for the conformal factor expressed in log-cylindrical coordinates. For the spherical and Euclidean gauge, we also set
$\mathcal{H}^{*}_{s}(u)=(\mathcal{P}_{s}\circ\mathfrak{F}^{-1})(u)\quad{\rm and}\quad\mathcal{H}_{s}(U)=(\mathcal{H}^{*}_{s}\circ\mathfrak{F}^{-1})(u\circ\Pi),$
where $v\in\mathcal{C}^{\infty}(\mathbf{R})$, $u\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$, and $U\in\mathcal{C}^{\infty}(\mathbf{S}^{n}\setminus\{N,S\})$ are solutions to ($\widehat{\mathcal{Q}}_{n,s}$), ($\overline{\mathcal{Q}}_{n,s}$), and (${\mathcal{Q}}^{\circ}_{n,s}$), respectively. Here we recall that $\Pi:\mathbf{S}\setminus\{N,S\}\to\mathbf{R}^{n}\setminus\{0\}$ is the stereographic projection and $\mathfrak{F}:\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})\to\mathcal{C}^{\infty}(\mathbf{R}\times\mathbf{S}^{n-1})$ is the log-cylindrical transform.
@@
Let $\{g_{k}=U_{k}^{\frac{4}{n-2s}}g_{\circ}\}$ be a sequence of singular Yamabe metrics such that all the asymptotic radial Pohozaev invariants are bounded away from zero, and the distances between singular points are also bounded away from zero. Using (for instance) [2, Lemma 11], it suffices to assume the singular set is a fixed finite set $\Lambda$ for the entire sequence. Now, let $\{D_{\ell}\}_{\ell\in\mathbf{N}}$ be a compact exhaustion of $\mathbf{S}^{n}\backslash\Lambda$ and for any large $\ell\gg 1$, we extract a convergent subsequence $U_{k}\rightarrow U_{*}(\ell)$. Letting $\ell\rightarrow\infty$, we obtain a limit $U_{*}\in\mathcal{C}^{2,\alpha}(\mathbf{S}^{n}\backslash\Lambda)$ for some $\alpha\in(0,1)$. Finally, we use the fact that the Pohozaev invariants are bounded away from zero to show that $U_{*}>0$ is strictly positive and defines a complete metric on $\mathbf{S}^{n}\backslash\Lambda$, i.e. $\lim_{p\to\Lambda}U(p)=\infty$. If $U_{*}\equiv 0$, we could rescale the sequence and contradict the positive lower bound for all Pohozaev invariants.
@@
In the regime in which we’ve established our classification, namely $1-\delta<s<1$, we expect the proof described above to carry over from the local setting to the nonlocal setting without any complications. It is worth remarking that in the local setting one can write the radial Pohozaev invariant very explicitly in terms of the necksize, while such a relation in the nonlocal setting is more elusive, mostly because the expression (45) is much more complicated. Thus, it would be very interesting to relate the necksize more explicitly to the radial Pohozaev invariant in the nonlocal setting.